# Continuous and Reversible Electrical Tuning of Fluorescent Decay Rate via Fano Resonance

### v1

@@ Abstract
Decay rate of an atomic or molecular dipole depends on the local density of optical states (LDOS) —Purcell effect—which can significantly be enhanced near a dielectric-core metal-shell nanoparticle (CSNP). On top of that, an auxiliary quantum object (QO) can introduce a Fano transparency in the plasmonic spectrum of the CSNP. Here, we show that an auxiliary QO, located at the hotspot of the CSNP, can modify the LDOS, hence the decay rate of an excited dipole. Moreover, by controlling the resonance of the auxiliary QO via an applied voltage, one can continuously and reversibly tune both the radiative and nonradiative decay rates of the dipole up to 2 orders-of-magnitude. This phenomenon emerges as an invaluable tool to implement in integrated quantum technologies, enabling realization of on-demand entanglement/single-photon sources, controlled execution of quantum gates and electrical-control of superradiant-like phase transitions. It also bears potential for application in super-resolution microscopy and surface-enhanced Raman spectroscopy (SERS).
@@ I Introduction
Therefore, achieving continuous and reversible control within a compact device geometry that provides large modulation depths with (and over) CPU clock speed compatibility remains as a great challenge in active control of fluorescence for technology applications.
@@
In this paper, we provide a proof-of-principle demonstration of active and reversible tuning of both radiative and nonradiative decay rates of a fluorescent molecule (FM). We electrically tune the LDOS by shifting the position of a Fano transparency (also known as Fano resonance) [23, 24, 25]. The resonance of the Fano transparency, induced by a quantum object (QO) residing at the CSNP hotspot (Fig. 1a), is tuned by shifting the level-spacing of the QO via an applied voltage. This way, the decay rates can be continuously tuned between =1–215 within a “picosecond” response time which is limited by the readout rate of the circuit operation (GB/s). We emphasize that the demonstrated system is fundamentally different from previously reported active PL switching and modulation effects that are based on screening of the PL for an external observer. By trapping the fluorescent state in the molecule (i.e., not quenching it), the nature of our switching process exhibits a crucial property with the ability to hold and release the interested fluorescent state by voltage application.
@@
In our demonstrated system, the presence of a plasmonic SiO2/Ag core (SiO2) shell (Ag) nanoparticle (CSNP) boosts the LDOS of a FM (Fig. 1a). When an auxiliary QO, such as a quantum dot (QD) or a defect-center of resonance is positioned near the CSNP, it introduces a Fano transparency at [26]. Fano transparency is the plasmon-analog of electromagnetically-induced-transparency (EIT) [27]. Coupling of a dark mode or a QO to a bright plasmon mode introduces two paths for the absorption [28]. When the two paths work out of phase, absorption is canceled, and a transparency window appears in the spectrum of the MNS [29]. Therefore, an absorption and scattering dip appears at . In this work, by choosing about the transition frequency of the FM’s metastable (fluorescence) state (Fig. 1b), we show that the presence of the QO turns off the LDOS that used to be enhanced due to the presence of the CSNP (Fig. 2a). In other words, at turns off the extra LDOS occurring due to the presence of the CSNP. Additionally, it also turns off the nonradiative decay rate emerging due to the dissipation at the Ag shell (Fig. 2b).
@@ II Results
Thus, we numerically demonstrate that we can actively control both the radiative and nonradiative decay rates of a FM for a narrow wavelength regime in a cavity-free environment. Our simulations were originally performed in vacuum therefore the dielectric environment of any possible device scheme (as shown in Figure 3a) would shift the plasmon resonances. This, however, won’t change the nature of the demonstrated phenomenon.
@@ III Conclusion, Discussion and Outlook
We provide a proof-of-principle numerical demonstration for the voltage-tunability of decay rates. Our simulations —exact solutions of Maxwell equations— show 2-orders magnitude modulation in the decay rate. A QO turns off the CSNP induced LDOS at fluorescence = . When is shifted by an applied voltage (Fig. 3a), the schemes (Figs. 2a and 2b) for the decay rates also shift, enabling a continuous tuning of these rates in Figs. 3b and 3c.
The phenomenon addresses a wide range of applications as we are demonstrating a fundamental issue in light-matter interaction, that is electrical-tuning of LDOS. This provides an invaluable tool especially for integrated quantum technologies.The technique offers ultimate control over on-demand single-photon sources and entanglement generation in circuits. These also imply voltage-controlled quantum gate operations. The phenomenon of superradiant phase transition, which depends on the decay rate, now can be electrically turned on/off in an integrated circuit. Thus, normal and superradiant phase transitions between zero and a very strong entanglement state [43], can be employed to serve as a controllable collective entanglement resource in the integrated circuits. In addition, quantum batteries can become voltage-controlled based on the demonstrated phenomenon [44, 45].
@@
We anticipate that the demonstrated technique can also be utilized in quantum optics for the on-demand generation of single-photons (or photon-pairs) and their synchronization with external systems such as a vibrating scanning probe microscopy (SPM) tip, or a confocal scanning laser microscope. Thanks to the ultimate control on the time gated activation of the emitter, the technique can also prove to be useful in performing spin-echo based quantum measurements/sensing applications such as quantum magnetometers [48] or quantum information processing [49]. Alternately, it can also serve as a voltage tunable band rejection filter for the radiative and nonradiative wavelengths falling into the corresponding Fano resonance linewidth. In general, the suitable execution of the demonstrated technique can form a base technology in the applications that require fine tuning of a specific fluorescence spectrum such as SERS and super-resolution microscopy.

### v2

# Continuous and reversible electrical-tuning of fluorescent decay rate via Fano resonance

@@ Abstract
We demonstrate that the decay rates of a fluorescent molecule can be controlled by electrically shifting a transparency introduced by a Fano resonance. An auxiliary quantum object (QO), located at the hotspot of a plasmonic nanoparticle, suppresses plasmonic excitation at its level spacing . As a result, the local density of states (LDOS) associated with the plasmonic spectrum is also suppressed at . By shifting via an applied voltage, we continuously tune the radiative and nonradiative decay rates of the fluorescent molecule by up to two orders of magnitude. This mechanism offers a valuable tool for integrated quantum technologies, enabling on-demand entanglement and single-photon sources, voltage-controlled quantum gate operations, and electrical control of superradiant-like phase transitions. The approach also holds promise for applications in super-resolution microscopy and surface-enhanced Raman spectroscopy (SERS).
@@ I Introduction
Therefore, achieving continuous and reversible control within a compact device geometry that provides large modulation depths with (and over) CPU clock speed compatibility remains as a great challenge in active control of fluorescence for technology applications. While certain progresses have been achieved on the electrical control of the Purcell effect [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], they either remain below the CPU clock speeds or possess poor modulation depths.
In this paper, we introduce a new method for fast electrical tuning of the Purcell effect that also achieves large modulation depths. Unlike conventional methods used in the literature [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], our approach relies on tuning a transparency in the plasmonic response that is introduced via a Fano resonance [23, 24, 25]. Our method outperforms existing electric-control techniques by offering a modulation depth of approximately 200 (i.e., 20,000%) and a response time on the order of a picosecond. It allows continuous tuning of the radiative and nonradiative decay rates of a fluorescent molecule, or the control of spontaneous parametric down-conversion in quantum integrated circuits.
An auxiliary quantum object (QO) —such as a quantum dot or a collection of defect centers [26, 27, 28]— is placed at the hotspot of a core-shell nanoparticle (CSNP) and introduces the Fano resonance (i.e., a transparency) [23, 24, 25], as shown in Fig. 1. This transparency appears at the level-spacing of the QO, which is = 766 nm in our simulations 111We refer the frequencies by the corresponding free-space wavelength, governed by . . The transparency suppresses plasmonic excitation and, in turn, eliminates the plasmonic enhancement in the local density of states (LDOS) at (see Fig. 2). When the metastable transition frequency of a fluorescent molecule —located 15 nm away— matches the QO level spacing (i.e., ), the plasmonic enhancement of fluorescence is reset to its vacuum value, (compare solid and dashed lines in Fig. 2).
The QO level spacing—and hence the spectral position of the transparency—can be controlled by an applied voltage (see Fig. 3a). For example, if shifts to = 767 nm, the new transparency appears at 767 nm. As a result, the LDOS associated with the decay at nm changes from to . By continuously shifting the QO level spacing, the decay rate can be tuned smoothly (see Fig. 3b and 3c). This method allows full control over the range to by adjusting the voltage from to 1 V. This mechanism enables our approach to achieve both large modulation depths and fast response times, limited only by the circuit readout rate (GB/s).
@@
We emphasize that the demonstrated system is fundamentally different from previously reported active PL switching and modulation effects that are based on screening of the PL for an external observer. By trapping the fluorescent state in the molecule (i.e., not quenching it), the nature of our switching process exhibits a crucial property with the ability to hold and release the interested fluorescent state by voltage application.
@@
We provide a proof-of-principle demonstration of active and reversible tuning of both radiative and nonradiative decay rates of a fluorescent molecule (FM) via exact solutions of the 3D Maxwell equations. We perform our simulations using in Lumerical –a reliable FDTD simulation toolbox [30]. In our demonstrated system, the presence of a plasmonic SiO2/Ag core (SiO2) shell (Ag) nanoparticle (CSNP) boosts the LDOS of a FM (Fig. 1a). When an auxiliary QO, such as a quantum dot (QD) or a defect-center of resonance , is positioned near the CSNP, it introduces a Fano transparency at [31]. Fano transparency is the plasmon-analog of electromagnetically-induced-transparency (EIT) [32]. Coupling of a dark mode or a QO to a bright plasmon mode introduces two paths for the absorption [33]. When the two paths work out of phase, absorption is canceled, and a transparency window appears in the spectrum of the MNS [34]. Therefore, an absorption and scattering dip appears at . In Sec. I of the Supplementary Material (SM) [35], we explain why and how such a transparency arises in the plasmonic excitation spectrum, using a single equation with detailed discussion.
In this work, by choosing about the transition frequency of the FM’s metastable (fluorescence) state (Fig. 1b), we show that the presence of the QO turns off the LDOS that used to be enhanced due to the presence of the CSNP (Fig. 2a). In other words, at turns off the extra LDOS occurring due to the presence of the CSNP. Additionally, it also turns off the nonradiative decay rate emerging due to the dissipation at the Ag shell (Fig. 2b).
@@ II Results
Thus, we numerically demonstrate that we can actively control both the radiative and nonradiative decay rates of a FM for a narrow wavelength regime in a cavity-free environment. Our simulations were originally performed in vacuum therefore the dielectric environment of any possible device scheme (as shown in Figure 3a) would shift the plasmon resonances. This, however, won’t change the nature of the demonstrated phenomenon. Moreover, in the supplementary material (SM) we demonstrate that such a control mechanism works for different choices of the FM-MNP distances (see Sec. IV of the Supplementary Material (SM) [35]).
@@ III Conclusion, Discussion and Outlook
We present a proof-of-principle numerical demonstration of voltage-tunable decay rates. Unlike conventional methods [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], we utilize a transparency introduced by a Fano resonance. This transparency occurs at the level spacing of an auxiliary quantum object (QO) placed at the hotspot of a plasmonic nanoparticle (see Fig. 1). We control the decay rate of a fluorescent molecule located 15 nm away from the nanoparticle. The transparency suppresses the LDOS induced by the nanoparticle (CSNP), which in turn reduces the decay rate of the fluorescent molecule. By shifting via an applied voltage, we control the position of the transparency and, consequently, the decay rate continuously (see Figs. 3b and 3c). Our simulations—based on exact solutions of the 3D Maxwell equations—demonstrate up to two orders of magnitude modulation in the decay rate.
The phenomenon addresses a wide range of applications as we are demonstrating a fundamental issue in light-matter interaction, that is electrical-tuning of LDOS. This provides an invaluable tool especially for integrated quantum technologies.The technique offers ultimate control over on-demand single-photon sources and entanglement generation in circuits. The method can also be utilized to control spontaneous parametric down conversion rate by choosing at . These also imply voltage-controlled quantum gate operations. The phenomenon of superradiant phase transition, which depends on the decay rate, now can be electrically turned on/off in an integrated circuit. Thus, normal and superradiant phase transitions between zero and a very strong entanglement state [45], can be employed to serve as a controllable collective entanglement resource in the integrated circuits. In addition, quantum batteries can become voltage-controlled based on the demonstrated phenomenon [46, 47].
@@
We anticipate that the demonstrated technique can also be utilized in quantum optics for the on-demand generation of single-photons (or photon-pairs) and their synchronization with external systems such as a vibrating scanning probe microscopy (SPM) tip, or a confocal scanning laser microscope. Unlike electrically-triggered quantum dots or defects [50], our method tunes the mean decay time after electrical excitation, and it can also be combined with such sources. Thanks to the ultimate control on the time gated activation of the emitter, the technique can also prove to be useful in performing spin-echo based quantum measurements/sensing applications such as quantum magnetometers [51] or quantum information processing [52]. Alternately, it can also serve as a voltage tunable band rejection filter for the radiative and nonradiative wavelengths falling into the corresponding Fano resonance linewidth. In general, the suitable execution of the demonstrated technique can form a base technology in the applications that require fine tuning of a specific fluorescence spectrum such as SERS and super-resolution microscopy.
@@ Acknowledgements.
- Note [1] We refer the frequencies by the corresponding free-space wavelength, governed by .
@@
- Yuan et al. [2002] Z. Yuan, B. E. Kardynal, R. M. Stevenson, A. J. Shields, C. J. Lobo, K. Cooper, N. S. Beattie, D. A. Ritchie, and M. Pepper, Electrically driven single-photon source, science 295, 102 (2002).